# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 4.3. Proof of Theorem B (iii)
Let $L$ be a regular leaf of the submetry $\pi\colon X\to Y$. Since $X$ has dimension $n$ and $Y$ is $k$-dimensional, $L$ has dimension $n-k$.
We start by proving that, under our hypotheses, the connected components of each leaf of the submetry are isometric to round spheres or round projective spaces.
As in the proof of Lemma 4.2, consider the group Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$, set $G=\operatorname{Ker}(\Psi)$, and let $\bar{x}\in X$ such that $\Sigma_{\bar{x}}(X)$ is isometric to the unit round $S^{n-1}$ and $\Sigma_{\pi(\bar{x})}(Y)$ is isometric to a unit round $S^{k-1}$. Each connected component $G(\bar{x})$ of $L_{\bar{x}}$ equipped with the intrinsic metric is a homogeneous space, locally compact, locally contractible, and since it is compact, each $1$-parameter subgroup of $G$ is rectifiable. Then, by [6, Theorem 7 I], $G(\bar{x})$ is a Finsler manifold. Moreover, since the spaces of directions of $L_{\bar{x}}$ are round $S^{n-k}$, the Minkowski cone of $G(\bar{x})$ is the Euclidean space $\mathbb{R}^{n-k}$. This implies that $G(\bar{x})$ with the intrinsic metric is a homogeneous Riemannian manifold. Therefore, $L_{\bar{x}}$ is a union of Riemannian manifolds and, hence, a compact Riemannian manifold with finitely many connected components. Moreover, by hypothesis, the isometry group of $L_{\bar{x}}$ has maximal dimension $(n-k)(n-k+1)/2$. By [28, Section II, Theorem 3.1], each connected component of $L_{\bar{x}}$ is isometric to a round $S^{n-k}$ or a round $\mathbb{R}P^{n-k}$. By Lemma 4.2, all connected components of $L_{\bar{x}}$ are homoeomorphic. Therefore, $L_{\bar{x}}$ is a finite union of either round spheres or round projective spaces.
Let us now show that $X$ is a topological manifold. Since every leaf of $\pi$ is a finite disjoint union of smooth manifolds and every point $x\in X$ is regular, we conclude that for any $x\in X$, the space of directions $\Sigma_{x}(X)$ is isometric to $V_{x}\ast H_{x}$. Recall that $V_{x}$ is the space of directions of $L_{x}$, and thus it is isomorphic to a unit round sphere $S^{n-k-1}$. The space $H_{x}$ is isometric to the space of directions $\Sigma_{\pi(x)}(Y)$, which is isometric to a unit round $S^{k-1}$. Therefore, the space of directions $\Sigma_{x}(X)$ is isometric to $S^{k-1}\ast S^{n-k-1}=S^{n-1}$. Hence, $X$ is a topological manifold.
We now show that $\pi\colon X\to Y$ is a fiber bundle. Note that, by [4], the map $\pi\colon X\to Y$ is in general at most a $C^{1,1}$ Riemannian submersion.
Recall taht $Y$ is connected. Fix $y_{0}\in Y$. Since $Y$ is a complete homogeneous Riemannian manifold, the injectivity radius is constant on $Y$. Fix $\varepsilon>0$ smaller than the injectivity radius, and consider the open cover $\{U_{\alpha}\}_{\alpha\in I}$ of $Y$ by normal balls of radius $\varepsilon$ centered at points $y_{\alpha}\in Y$. When $Y$ is a round sphere $S^{k}$ or a round projective space $\mathbb{R}P^{k}$, we avoid choosing $y_{\alpha}$ to be the conjugate point to $y_{0}$. For each $\alpha\in I$, let $c_{\alpha}\colon[0,1]\to Y$ be the unique minimizing geodesic joining $y_{0}$ to $y_{\alpha}$, set $F=\pi^{-1}(y_{0})\subset X$, and let $H=\operatorname{\mathrm{Homeo}}(F,F)=\{h\colon F\to F\mid\text{$h$ is a homeomorphism}\}$.
For $y\in Y$, set $F_{y}=\pi^{-1}(y)\subset X$ and let
$G_{y}=\operatorname{\mathrm{Homeo}}(F,F_{y})=\{h\colon F\to F_{y}\mid\text{$h$is a homeomorphism}\}.$
Let $P=\sqcup_{y\in Y}G_{y}$ and consider the map $\Upsilon\colon P\to Y$ given by $\Upsilon(g)=y$ if $g\in G_{y}$. We define an action of $H$ on $P$ as follows: given $h\in H$ and $g\in G_{y}\subset P$, we set $h\cdot g=g\circ h^{-1}\colon F\to F_{y}$, which is an element of $G_{y}$.
For each $y\in U_{\alpha}$, there is a unique minimizing geodesic $c_{y,\alpha}\colon[0,1]\to U_{\alpha}$ joining $y_{\alpha}$ to $y$. This geodesic is completely determined by its initial tangent vector $c_{y,\alpha}^{\prime}(0)\in T_{y_{\alpha}}Y$. Moreover, since $C(\Sigma_{x}(X))=C(H_{x})\times C(V_{x})$, for each unique minimizing geodesic $c\colon I\to Y$ in $Y$, given $x\in\pi^{-1}(c(0))$, there is a unique geodesic lift $\tilde{c}_{x}\colon I\to X$ by Lemma 2.14. Thus, we have a holonomy homeomorphism $h_{c}\colon\pi^{-1}(c(0))\to\pi^{-1}(c(1))$ given by $h_{c}(x)=\tilde{c}_{x}(1)$.
Define $s_{\alpha}\colon U_{\alpha}\to\pi^{-1}(U_{\alpha})$ by setting
$s_{\alpha}(y)=h_{c_{y,\alpha}}\circ h_{c_{\alpha}}\colon F\to F_{y}.$
Observe that this is a local section for the projection $\pi\colon P\to Y$ by construction; i.e., $\pi(s_{\alpha}(y))=y$ for all $y\in Y$.
Let us now show that the map $\Phi_{\alpha}\colon U_{\alpha}\times H\to\pi^{-1}(U_{\alpha})$ defined as
$\Phi_{\alpha}(y,h)=s_{\alpha}(y)\circ h^{-1}\colon F\to F_{y}$
is a bijection. Suppose that $\Phi_{\alpha}(y,h)=\Phi_{\alpha}(y^{\prime},h^{\prime})$. Then $y=\Upsilon(\Phi_{\alpha}(y,h))=\Upsilon(\Phi_{\alpha}(y^{\prime},h^{\prime}))=y^{\prime}$. Thus, $h^{-1}=s_{\alpha}(y)^{-1}\circ s_{\alpha}(y)\circ(h^{\prime})^{-1}=(h^{\prime})^{-1}$, which implies that $\Phi_{\alpha}$ is injective. Given a homeomorphism $g\colon F\to F_{y}$, consider $s_{\alpha}(y)\colon F\to F_{y}$. Then $h=g^{-1}\circ s_{\alpha}(y)$ is an element in $G$. Thus, $\Phi_{\alpha}(y,h)=s_{\alpha}(y)\circ h=g$, which implies that $\Phi_{\alpha}$ is onto. Thus, we may induce a topology on $\Upsilon^{-1}(U_{\alpha})$ making $\Phi_{\alpha}$ into a homeomorphism. Moreover, we have computed the inverse of $\Phi_{\alpha}$, which is given by
$\Phi_{\alpha}^{-1}(g)=(\Upsilon(g),g^{-1}\circ s_{\alpha}(\Upsilon(h))).$
For $U_{\alpha}\cap U_{\beta}\neq\emptyset$ we have:
$\Phi_{\beta}^{-1}\circ\Phi_{\alpha}(y,h)=\Phi_{\beta}^{-1}(s_{\alpha}(y)\circ h^{-1}))=(y,h\circ s_{\alpha}(y)^{-1}\circ s_{\beta}(y)).$
Thus, transition functions are of the form $t_{\alpha\beta}=s_{\alpha}^{-1}\circ s_{\beta}\colon U_{\alpha}\cap U_{\beta}\to H$ and are continuous by construction, and satisfy the cocycle condition.
Therefore, $\Upsilon\colon P\to Y$ is a continuous principal bundle. We see that $X$ is the associated fiber bundle via the homeomorphism $\varphi\colon P\times_{H}F\to X$ given by $\varphi([g,v])=g(v)$. ∎

### v3

@@ 4.4 Disconnected fiber case
By the factorization of submetries between Alexandrov spaces [36, Theorem 10.1], there exist submetries $\tilde{\pi}\colon X\to Z$ with connected fibers and $\pi_{D}\colon Z\to Y$ with discrete fibers such that $\pi=\pi_{D}\circ\tilde{\pi}$}.
[…]
$\pi^{-1}(y)=\bigsqcup_{z\in\pi^{-1}_{D}(y)}\tilde{\pi}^{-1}(z)$
[…]
Since $\dim(Z)=\dim(Y)=k$, we have
[…]
$\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2},$
[…]
By Lemma 4.5, each fiber of $\tilde{\mathcal{F}}$ (with the intrinsic metric) is isometric to a round sphere or a round real projective space. Therefore, each fiber of $\pi$ is a disjoint union of round spheres or real projective spaces. Since $\pi$ is proper, each fiber $\pi^{-1}(y)$ is compact, hence consists of finitely many connected components.
By Lemma 4.7, $\tilde{\pi}\colon X\to Z$ is a smooth Riemannian submersion. Since $\pi_{D}\colon Z\to Y$ has discrete fibers, $\pi_{D}$ is a (Riemannian) covering map (see [34, Theorem 1.2]). Since $Y$ is simply-connected, the covering $\pi_{D}\colon Z\to Y$ is trivial. Thus, $Z=Y$ and $\pi_{D}$ is an isometry. Hence, $\pi=\tilde{\pi}$ is a smooth Riemannian submersion with connected fibers (see also [39]). In particular, $\pi\colon X\to Y$ is a fiber bundle with fiber $F$ a single round sphere or real projective space. Since $Y$ is contractible, the bundle is trivial and $X$ must be homeomorphic to $Y\times F$. ∎
Lemma 4.9.
If the fibers of $\pi\colon X\to Y$ are possibly disconnected and $Y$ is $S^{k}$ or $\mathbb{R}P^{k}$, then for $k\geq 2$} either each fiber has exactly two connected components and $Y$ is isometric to a round $\mathbb{R}P^{k}$, or each fiber is connected and $Y$ is isometric to a round $S^{k}$ or a round $\mathbb{R}P^{k}$. For k=1{k=1}, the fibers can have $m\geq 1$ connected components.
Proof.
Factor $\pi=\pi_{D}\circ\tilde{\pi}$ with $\tilde{\pi}\colon X\to Z$ a submetry with connected fibers, and $\pi_{D}\colon Z\to Y$ a submetry with discrete fibers. Arguing as in the proof of Lemma 4.8, each fiber of $\tilde{\pi}$ is isometric to a round sphere or a round real projective space and $\pi_{D}\colon Z\to Y$ is a Riemannian covering map with $Y$ isometric to a round $S^{k}$ or a round $\mathbb{R}P^{k}$. As $Y$ is connected, the number of sheets of the covering $\pi_{D}$ is constant. Therefore, the number of connected components of the fibers of $\pi$ is constant on $Y$.
If $k\geq 2$, then either $Z=Y=S^{k}$ and $\pi_{D}$ is an isometry, or $Z=S^{k}$ and $Y=\mathbb{R}P^{k}$ with $\pi_{D}$ the standard two-fold covering. If $\pi_{D}$ is an isometry, then $\pi=\tilde{\pi}$ and the fibers are connected. Hence, $\pi=\tilde{\pi}$ is a smooth Riemannian submersion. In particular, $\pi\colon X\to Y$ is a fiber bundle, with fiber a round sphere or real projective space and base a sphere. If $\pi_{D}$ is the two-fold cover $S^{k}\to\mathbb{R}P^{k}$, then for each $y\in Y$, $\pi^{-1}_{D}$ consists of two points and $\pi^{-1}(y)$ has exactly two connected components.
If $k=1$, in addition to the identity $S^{1}\to S^{1}$ and the two-fold covering $S^{1}\to\mathbb{R}P^{1}$, one must also consider the standard $m$-fold Riemannian coverings $S^{1}\to S^{1}$ given by cyclic rotation groups of order $m$. ∎
Acknowledgements
We thank Alexander Lytchak and Marco Radeschi for helpful comments on a preliminary version of this article. We thank the organizers of the IV joint meeting of RSME and SMM, and the Universidad Politécnica de Valencia, for their hospitality while this manuscript was finished. We thank the anonymous referees for suggestions that improved both clarity and accuracy. In particular, we are grateful to one of the referees for observations that helped strengthen Theorems A and B and simplify their proofs, for pointing out the argument in the proof of Proposition 2.23, and for proposing the question highlighted in Remark 1.5. D. Corro was supported in part by UNAM-DGAPA Postdoctoral fellowship of the Institute of Mathematics, and by the DFG (grant CO 2359/1-1, Priority Programme SPP2026 “Geometry at Infinity”), and by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey]. F. Galaz-García was supported in part by the DFG (grant GA 2050 2-1, Priority Programme SPP2026 “Geometry at Infinity”).